# Streaming Federated Learning with Markovian Data

### v1

@@ 4.1 Assumptions
To conduct the convergence analysis, we first impose the following assumptions on the objective functions and the stochastic gradients, which allow for non-convexity.
@@
$F(w_{2})$ $\leq F(w_{1})+\langle\nabla F(w_{1}),w_{2}-w_{1}\rangle$ (7)
$+\frac{L}{2}\left\lVert w_{1}-w_{2}\right\rVert^{2}.$ (8)
@@
$f_{m}(w_{2};x)$ $\leq f_{m}(w_{1};x)+\langle\nabla f_{m}(w_{1};x),w_{2}-w_{1}\rangle$ (9)
$+\frac{L}{2}\left\lVert w_{1}-w_{2}\right\rVert^{2}.$ (10)
@@
Assumption 4.3 uniformly bounds the gradient estimation error for each client. While stronger than the bounded variance assumption, this assumption is standard for stochastic optimization with Markovian noise in both the centralized [Beznosikov et al., 2023] and federated [Mangold et al., 2024] settings.
Heterogeneity in client objective functions and data distributions is a well-known difficulty in FL algorithms based on local SGD, which leads to client drift [Karimireddy et al., 2020]. In order to provide convergence guarantees in the i.i.d. setting, it is necessary to impose constraints on the local gradient norms [Woodworth et al., 2020, Yun et al., 2022]. Client heterogeneity poses the same difficulty in the Markovian setting. As such, we impose the following bounded gradient dissimilarity (BGD) assumption.
@@
This assumption was first introduced in [Karimireddy et al., 2020] and has since become ubiquitous in the analysis of local SGD [Yun et al., 2022]. In Section 4, we establish iteration and communication complexity bounds of local SGD under these assumptions.
Finally, we define the following three classes of problems for the analysis of minibatch SGD, local SGD, and local SGD-M. Recall from Sec. 3.2 that the system-level Markov chain is denoted by $X$. We assume throughout that all objective functions are in general non-convex and $F$ is bounded from below by $F^{*}>-\infty$.
[…]
@@
$\mathcal{F}_{3}(L,\sigma,\theta,\delta,\tau)\subset\mathcal{F}_{2}(L,\sigma,\tau)\subset\mathcal{F}_{1}(L,\sigma,\tau).$ (19)
In the following sections, we analyze the convergence of Minibatch SGD, Local SGD, and Local SGD-M for the problem classes $\mathcal{F}_{1},\mathcal{F}_{3}$, and $\mathcal{F}_{2}$, respectively. We establish conditions, for any $\epsilon>0$, on the step sizes, the number of local steps $K$, and the number of communication rounds $T$ in order to guarantee
@@ 4.2 Minibatch SGD
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}&\left(\frac{\Delta_{0}}{\gamma T}+\frac{\tau\sigma^{2}}{MK}\right.\\ &\left.+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}\right).\end{split}$ (21)
The proof of Theorem 4.5 is given in Appendix D. The first deterministic term depends on the initial iterate and decays with the same rate as in the i.i.d setting. The last two terms arise from the dependence structure and non-stationarity of the Markov data process. In contrast to i.i.d. sampling [Woodworth et al., 2020], these terms cannot be controlled by the step size. In particular, the second term reveals that the stochastic gradient noise is amplified by the mixing time of the system-level Markov chain, and the last term is due to the bias of Markovian sampling whenever the system-level Markov chain is non-stationary.
@@
  1. (i)
    If the system-level Markov chain $X$ is stationary,
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\quad T\geq\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\right).$ (22)
  2. (ii)
    If the system-level Markov chain $X$ is non-stationary, with $\epsilon\leq\mathcal{O}\left(\sqrt{\frac{\sigma^{2}}{M}}\right)$,
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\quad T\geq\mathcal{O}\left(\frac{L\Delta_{0}\vee\sigma^{2}}{\epsilon^{2}}\right).$ (23)
The proof of Corollary 4.6 follows directly from Theorem 4.5 and is given in Appendix D.
@@ 4.3 Local SGD
In the analysis of Local SGD, bounded heterogeneity assumptions are required, even in the i.i.d. setting [Karimireddy et al., 2020]. We consider the problem class $\mathcal{F}_{3}(L,\sigma,\theta,\delta,\tau)$, which requires sample-wise smoothness (Assumption 4.2) and bounded heterogeneity (Assumption 4.4).
@@
For any $F\in\mathcal{F}_{3}(L,\sigma,\theta,\delta,\tau)$, with local step size $\eta\leq\mathcal{O}\left(\frac{1}{LK\delta^{2}}\right)$, the iterates of Local SGD satisfy:
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}&\left(\frac{\Delta_{0}}{\eta KT}+\frac{\tau\sigma^{2}}{MK}\right.\\ &\left.+\frac{LK\eta\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}\right.\\ &\left.+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert\mathcal{Q}P^{Kt}-\pi\right\rVert_{TV}\right).\end{split}$ (24)
The proof of Theorem 4.7 is given in Appendix E. The convergence behavior of Local SGD closely resembles that of Minibatch SGD, except for the additional third term introduced by the client drift. We note that this term can be controlled by having a sufficiently small step size. Similar to Corollary 4.6, the following corollary specifies the communication and sample complexity of Local SGD for both stationary and non-stationary Markov chains.
@@
  1. (i)
    If the system-level Markov chain $X$ is stationary,
    $K$ $\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),$ (26)
    $T$ $\geq\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right).$ (27)
  2. (ii)
    If the system-level Markov chain $X$ is non-stationary, with $\epsilon\leq\mathcal{O}\left(\sqrt{\frac{\sigma^{2}}{M}}\right)$,
    $K$ $\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),$ (28)
    $T$ $\geq\mathcal{O}\left(\frac{L\Delta_{0}\vee\sigma^{2}}{\epsilon^{2}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right).$ (29)
The proof of Corollary 4.8 is given in Appendix E.
@@ 4.4 Local SGD with Momentum
In the case of i.i.d sampling, several algorithms have been proposed to mitigate the client drift effect without the need of Assumption 4.4. One line of work focuses on the control variate technique, first introduced by Karimireddy et al. [2020]. Recently, the work by Cheng et al. [2024], Yan et al. [2025] has demonstrated that momentum can also be used to mitigate client drift in the i.i.d. setting.
In this section, we study the convergence of a momentum-based variant of Local SGD, named Local SGD-M, detailed in Algorithm 1. The key difference from Local SGD is that the local updates are computed via the convex combination of the gradient estimate for the local objective and the aggregated updates from the previous communication round. We show that, with Markovian data, this technique also mitigates the impact of client drift. In particular, the lower bound on communication complexity in the i.i.d. setting [Patel et al., 2022] is achieved up to a constant without any heterogeneity assumptions.
@@
        Sample $(x_{t}^{m,k})_{k=1}^{K}$
        for $k=0$ to $K-1$ do
           $v_{t}^{m,k}=\beta\nabla f_{m}\left(w_{t}^{(m,k)};x_{t}^{(m,k)}\right)+\left(1-\beta\right)v_{t}$
@@
The convergence result for Local SGD with momentum holds for the broader function class $\mathcal{F}_{2}(L,\sigma,\tau)$, which retains the sample-wise smoothness condition but relaxes the requirement for bounded heterogeneity.
@@
$\begin{split}\eta KL\leq\mathcal{O}&\left(\frac{1}{\beta}\wedge\frac{1}{\sqrt{MK\beta^{2}}}\wedge\sqrt{\frac{L\Delta_{0}}{\beta^{3}TG_{0}}}\wedge\frac{1}{\left(1-\beta\right)}\right.\\ &\leavevmode\nobreak\ \leavevmode\nobreak\ \left.\wedge\frac{1}{\beta\gamma LT}\right),\\ \gamma\leq\mathcal{O}&\left(\frac{\beta}{L}\right),\end{split}$ (30)
@@
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}&\left(\frac{L\Delta_{0}}{\beta T}+\frac{\tau\sigma^{2}}{MK}\right.\\ &\left.+\frac{\sigma^{2}}{T}\sum_{t=0}^{T-1}\left\lVert QP^{Kt}-\pi\right\rVert_{TV}\right).\end{split}$ (31)
The proof of Theorem 4.9 is given in Appendix F.6. The communication and sample complexities of Local SGD-M are characterized in the following corollary.
@@
  1. (i)
    If the system-level Markov chain $X$ is stationary,
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\leavevmode\nobreak\ T\geq\mathcal{O}\left(\frac{L\Delta_{0}}{\beta\epsilon^{2}}\right).$ (32)
  2. (ii)
    If the system-level Markov chain $X$ is non-stationary, with $\epsilon\leq\mathcal{O}\left(\sqrt{\frac{\sigma^{2}}{M}}\right)$,
    $K\geq\mathcal{O}\left(\frac{\tau\sigma^{2}}{M\epsilon^{2}}\right),\leavevmode\nobreak\ T\geq\mathcal{O}\left(\frac{\frac{L\Delta_{0}}{\beta}\vee\sigma^{2}}{\epsilon^{2}}\right).$ (33)
The proof of Corollary 4.10 is given in Appendix F.7.
@@ 4.5 Discussion
Table 1 summarizes the communication and sample complexities of Minibatch and Local SGD, and the momentum-based variant of Local SGD, for both stationary and non-stationary Markovian data streams.
The communication complexity is measured by the total number of communication rounds, while the sample complexity refers to the total number of samples processed. For our FL algorithms, the sample complexity is equal to $KT$ and can be estimated by multiplying the bounds for $K$ and $T$ given in Corollaries 4.6, 4.8, and 4.10.
In the stationary case, at the cost of more local computation, Minibatch SGD can achieve the lower bound in Patel et al. [2022], while Local SGD suffers from client drift, and can only reach this lower bound in the low-heterogeneity low-noise regime; i.e., when $\frac{\theta^{2}+\sigma^{2}}{\epsilon^{2}}\leq\delta^{4}$.
In the case of a non-stationary system-level Markov chain, an additional term $\frac{\sigma^{2}}{\epsilon^{2}}$ is introduced in the communication complexity of both Minibatch and Local SGD. Convergence can only be established in the high-precision regime (small $\epsilon$). Minibatch SGD can only achieve the lower bound in Patel et al. [2022] if the gradient noise is sufficiently small $\sigma^{2}\leq\mathcal{O}(L\Delta_{0})$.
@@
Local SGD-M successfully mitigates the client drift caused by heterogeneity. In particular, the lower bound on the communication complexity in Patel et al. [2022] is achieved without any heterogeneity assumptions, up to a constant of $1/\beta$, while maintaining the same number of samples per communication round $K$. This contrasts with SCAFFLSA [Mangold et al., 2024], which utilizes control variates to mitigate heterogeneity, resulting in lower communication complexity. However, SCAFFLSA needs a higher number of samples per communication round than its vanilla version FedLSA and still relies on a heterogeneity assumption around the optimum.
[…]
Table 1: Comparison of Algorithms in Terms of Communication and Sample Complexity
Communication $T$ Sample complexity $KT$
Lower bound i.i.d. [Patel et al., 2022] $\Omega\left(\frac{L\Delta_{0}}{\epsilon^{2}}\right)$ $\Omega\left(\frac{L\Delta_{0}\sigma}{M\epsilon^{3}}\right)$
MAG [Dorfman and Levy, 2022]* $\tilde{\mathcal{O}}\left(\tau_{m}\frac{(M+L+G)^{2}G^{2}}{\epsilon^{4}}\right)$
Stationary Minibatch SGD $\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\right)$ $\mathcal{O}\left(\tau\frac{L\Delta_{0}\sigma^{2}}{M\epsilon^{4}}\right)$
Local SGD $\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right)$ $\mathcal{O}\left(\tau\frac{L\Delta_{0}\sigma^{2}}{M\epsilon^{4}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right)$
Local SGD-M $\mathcal{O}\left(\frac{L\Delta_{0}}{\beta\epsilon^{2}}\right)$ $\mathcal{O}\left(\tau\frac{L\Delta_{0}\sigma^{2}}{\beta M\epsilon^{4}}\right)$
Non-stationary Minibatch SGD $\mathcal{O}\left(\frac{L\Delta_{0}\vee\sigma^{2}}{\epsilon^{2}}\right)$ $\mathcal{O}\left(\tau\frac{\left(L\Delta_{0}\vee\sigma^{2}\right)\sigma^{2}}{M\epsilon^{4}}\right)$
Local SGD $\mathcal{O}\left(\frac{L\Delta_{0}\vee\sigma^{2}}{\epsilon^{2}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right)$ $\mathcal{O}\left(\tau\frac{\left(L\Delta_{0}\vee\sigma^{2}\right)\sigma^{2}}{M\epsilon^{4}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right)$
Local SGD-M $\mathcal{O}\left(\frac{\frac{L\Delta_{0}}{\beta}\vee\sigma^{2}}{\epsilon^{2}}\right)$ $\mathcal{O}\left(\tau\frac{\left(\frac{L\Delta_{0}}{\beta}\vee\sigma^{2}\right)\sigma^{2}}{M\epsilon^{4}}\right)$
    Here $\tau_{m}$ is the mixing time of client $m$’s Markov chain, $M$ and $G$ are upper bounds on the domain diameter and on the gradient norm, respectively, and $\tilde{\mathcal{O}}$ ignores logarithmic constants.

### v2

@@ 4.1 Assumptions
To conduct the convergence analysis, we first impose the following assumptions on the objective functions and the stochastic gradients.
@@
$F(w_{2})$ $\leq F(w_{1})+\langle\nabla F(w_{1}),w_{2}-w_{1}\rangle+\frac{L}{2}\left\lVert w_{1}-w_{2}\right\rVert^{2}.$ (4)
@@
$f_{m}(w_{2};x)$ $\leq f_{m}(w_{1};x)+\langle\nabla f_{m}(w_{1};x),w_{2}-w_{1}\rangle+\frac{L}{2}\left\lVert w_{1}-w_{2}\right\rVert^{2}.$ (5)
@@
Assumption 4.3 uniformly bounds the gradient estimation error for each client. While stronger than the bounded variance assumption, this assumption is standard for stochastic optimization with Markovian noise in both the centralized [3] and federated [37] settings.
Heterogeneity in client objective functions and data distributions is a well-known challenge for FL algorithms based on Local SGD, as it leads to client drift [27]. In order to provide convergence guarantees in the i.i.d. setting, it is necessary to impose constraints on the local gradient norms [60, 64]. Client heterogeneity poses the same difficulty in the Markovian setting. As such, we impose the following bounded gradient dissimilarity (BGD) assumption.
@@
This assumption was first introduced in [27] and has since become ubiquitous in the analysis of Local SGD [64]. We highlight that the BGD assumption includes the assumptions in [60, 33]. While weaker heterogeneity assumptions exist in the literature, these are not straightforward to apply for non-convex objectives where the BGD assumption is standard [30]. In Section 4, we establish iteration and communication complexity bounds of Local SGD under these assumptions.
Finally, we define the following three classes of problems for the analysis of Minibatch SGD, Local SGD, and Local SGD with Momentum (SGD-M). Recall from Section 3.2 that the system-level Markov chain is denoted by $X$. We assume throughout that all objective functions are in general non-convex and $F$ is bounded from below by $F^{*}>$-$\infty$.
@@
$\mathcal{F}_{3}(L,\sigma,\theta,\delta,\nu_{ps},C_{\infty})\subset\mathcal{F}_{2}(L,\sigma,\nu_{ps},C_{\infty})\subset\mathcal{F}_{1}(L,\sigma,\nu_{ps},C_{\infty}).$ (11)
In the following sections, we analyze the convergence of Minibatch SGD, Local SGD, and Local SGD-M for the problem classes $\mathcal{F}_{1},\mathcal{F}_{3}$, and $\mathcal{F}_{2}$, respectively. For any $\epsilon>0$, we derive conditions on the step sizes, the number of local steps $K$, and the number of communication rounds $T$ that ensure convergence to an $\epsilon$-accurate solution, i.e.,
@@ 4.2 Minibatch SGD
For the problem class $\mathcal{F}_{1}(L,\sigma,\nu_{ps},C_{\infty})$, with global step size $\gamma\leq 1/L$, the iterates of Minibatch SGD satisfy:
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}\left(\frac{\Delta_{0}}{\gamma T}+\frac{C_{\infty}\sigma^{2}}{\nu_{ps}MK}\right).\end{split}$ (13)
The proof of Theorem 4.5 is given in Appendix D.1. The first deterministic term depends on the initial iterate and decays with the same rate as in the i.i.d. setting. The last term arises from the dependence structure and non-stationarity of the Markov data process. In contrast to i.i.d. sampling [60], these terms cannot be controlled by the step size. In particular, Theorem 4.5 reveals that the stochastic gradient noise is amplified by a factor inversely proportional to the spectral gap of the system-level Markov chain.
@@
$K=\mathcal{O}\left(\frac{C_{\infty}\sigma^{2}}{\nu_{ps}M\epsilon^{2}}\right),\quad T=\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\right).$
The proof of Corollary 4.6 is deferred to Appendix D.1.
@@ 4.3 Local SGD
In the analysis of Local SGD, bounded heterogeneity assumptions are required, even in the i.i.d. setting [27]. We consider the problem class $\mathcal{F}_{3}(L,\sigma,\theta,\delta,\nu_{ps},C_{\infty})$, which requires sample-wise smoothness (Assumption 4.2) and bounded heterogeneity (Assumption 4.4).
@@
For any $F\in\mathcal{F}_{3}(L,\sigma,\theta,\delta,\nu_{ps},C_{\infty})$, with local step size $\eta\leq\mathcal{O}\left(\frac{1}{LK\delta^{2}}\right)$, the iterates of Local SGD satisfy:
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}&\left(\frac{\Delta_{0}}{\eta KT}+\frac{C_{\infty}\sigma^{2}}{\nu_{ps}MK}\right.\left.+\frac{LK\eta\left(\theta^{2}+\sigma^{2}\right)}{\delta^{2}}\right).\end{split}$ (14)
The proof of Theorem 4.7 is given in Appendix D.2. The convergence behavior of Local SGD closely resembles that of Minibatch SGD, except for the additional third term introduced by the client drift. We note that this term can be controlled by having a sufficiently small step size. Similar to Corollary 4.6, the following corollary specifies the communication and sample complexity of Local SGD for both stationary and non-stationary Markov chains.
@@
$K=\mathcal{O}\left(\frac{C_{\infty}\sigma^{2}}{\nu_{ps}M\epsilon^{2}}\right),\quad T=\mathcal{O}\left(\frac{L\Delta_{0}}{\epsilon^{2}}\left(\delta^{2}\vee\frac{\theta^{2}+\sigma^{2}}{\delta^{2}\epsilon^{2}}\right)\right).$
The proof of Corollary 4.8 is also given in Appendix D.2.
@@ 4.4 Local SGD with Momentum
In the case of i.i.d. sampling, several algorithms have been proposed to mitigate the client drift effect without the need of Assumption 4.4. One line of work focuses on the control variate technique, first introduced by [27]. Recently, [9, 61] have demonstrated that momentum can also be used to mitigate client drift in the i.i.d. setting.
In this section, we study the convergence of a momentum-based variant of Local SGD, named Local SGD-M, detailed in Algorithm 1. The key difference from Local SGD is that the local updates are computed via the convex combination of the gradient estimate for the local objective and the aggregated updates from the previous communication round. This allows us to build a recursive bound for the gradient estimate error, without relying on any heterogeneity assumption to bound the drift caused by local updates as in the case of Local SGD.
We show that, with Markovian data, this technique also mitigates the impact of client drift. In particular, the lower bound on communication complexity in the i.i.d. setting [42] is achieved up to a constant without any heterogeneity assumptions.
@@
   for $k=0$ to K1K-1 do
    $v_{t}^{(m,k)}=\beta\nabla f_{m}\left(w_{t}^{(m,k)};x_{t}^{(m,k)}\right)+\left(1-\beta\right)v_{t}$
@@
The convergence result for Local SGD with momentum holds for the broader function class $\mathcal{F}_{2}(L,\sigma,\nu_{ps},C_{\infty})$, which retains the sample-wise smoothness condition but relaxes the requirement for bounded heterogeneity.
@@
For the problem class $\mathcal{F}_{2}(L,\sigma,\nu_{ps},C_{\infty})$, with the following conditions on the step sizes:
$\begin{split}\eta KL\leq\mathcal{O}&\left(\frac{1}{\beta}\wedge\sqrt{\frac{C_{\infty}}{\nu_{ps}MK\beta^{2}}}\wedge\sqrt{\frac{L\Delta_{0}}{\beta^{3}TG_{0}}}\wedge\frac{1}{\left(1-\beta\right)}\right.\left.\wedge\frac{1}{\beta\gamma LT}\right),~~~\gamma\leq\mathcal{O}\left(\frac{\beta}{L}\right),\end{split}$ (16)
@@
$\begin{split}\mathbb{E}\left[\left\lVert\nabla F(\hat{w}_{T})\right\rVert^{2}\right]\leq\mathcal{O}&\left(\frac{L\Delta_{0}}{\beta T}+\frac{C_{\infty}\sigma^{2}}{\nu_{ps}MK}\right).\end{split}$ (17)
The proof of Theorem 4.9 is given in Appendix D.3. The communication and sample complexities of Local SGD-M are characterized in the following corollary.
@@
$K=\mathcal{O}\left(\frac{C_{\infty}\sigma^{2}}{\nu_{ps}M\epsilon^{2}}\right),\quad T=\mathcal{O}\left(\frac{L\Delta_{0}}{\beta\epsilon^{2}}\right).$
The proof of Corollary 4.10 is delegated to Appendix D.3.
@@ 4.5 Discussion
Impact of Markovian Data: Our convergence analysis in Corollaries 4.6, 4.8, and 4.10 reveals the impact of Markovian sampling for smooth non-convex objectives. Due to temporal dependence between samples, the gradient noise is scaled by $C_{\infty}/\nu_{ps}$, which depends on the spectral properties of the Markov kernel. Moreover, the stochastic gradient estimates computed by the clients are inherently biased. This bias arises from the Markovian structure of the data and, unlike the i.i.d. case, cannot be controlled by having a small or decaying step size. This is due to the fact that the bias contributes to the terms involving the gradient noise $\sigma^{2}$ in our bounds. Consequently, in order to control these terms and guarantee convergence, the number of local steps $K$ must be sufficiently large.
However, we emphasize that our results remain consistent with the core principles of FL: leveraging more local computation to reduce communication costs. Indeed, a comparison with [42, Table 2] shows that the lower bound established in the i.i.d. setting for communication complexity can be achieved by Minibatch SGD and Local SGD-M. Similar results for the special case of Federated Stochastic Approximation (FSA) can be observed in [37, Corollary 4.5], where additional local computation is required to reduce the third term in [37, Theorem 4.1].
Some prior works in the centralized [3, 15, 14] or the token-passing [17] setting introduce auxiliary hyper-parameters that can be tuned to make the bias vanish, instead of tuning directly the step size. This is achieved:
  • in [17, 14] by decomposing the bias into other terms involving delayed gradients, and scaling such delays as $1/\sqrt{T}$. However, this technique requires a constraint on the minimum number of iterations $T$ (analogous to the number of communication rounds in our analysis), at the order of the mixing time of the Markov chain. Such a constraint is undesirable in the FL setting, where the main goal is to minimize communication. A recent paper on FSA [65] follows the same strategy and indeed needs to impose a lower bound on $T$ proportional to the maximum mixing time across all clients.
  • in [3, 15] by using minibatch gradient estimates (instead of using just one sample). More precisely, at each iteration, a random batch size is drawn from a truncated geometric distribution with maximum value $m$, requiring a new hyper-parameter. [3, 15] show that the bias scales inversely with $m$, which is then chosen to be scaled with $\sqrt{T}$.
We highlight that the sample complexities obtained in our paper are of the order of $1/(M\epsilon^{4})$, matching those achieved in [3, 15, 14, 17], scaled inversely with the number of clients $M$. However, as already mentioned in Section 3.2, our assumptions on the underlying Markov processes are more general. Furthermore, as the terms involving $\sigma^{2}$ scale with $1/K$, following the approach in [3, 15] and letting $K$ scale with $T$ would indeed drive these terms to zero as $T\rightarrow\infty$. Nevertheless, doing so would not lead to any improvement in the overall sample complexity. Further details on this, as well as a more exhaustive discussion on the differences of our work with existing works, are provided in Appendix C.
@@
Impact of Heterogeneity: Convergence analysis of Local SGD requires the BGD heterogeneity assumption. The lower bound for communication complexity in the i.i.d. setting can only be achieved within the low-heterogeneity, low-noise regime; i.e., when $\frac{\theta^{2}+\sigma^{2}}{\epsilon^{2}}\leq\delta^{4}$. In contrast, Local SGD-M successfully mitigates the client drift caused by heterogeneity. In particular, the lower bound on the communication complexity in [42] is achieved without any heterogeneity assumptions, up to a constant of $1/\beta$, while maintaining the same number of samples per communication round $K$. This contrasts with SCAFFLSA [37], which utilizes control variates to mitigate heterogeneity, resulting in lower communication complexity. However, SCAFFLSA needs a higher number of samples per communication round than its vanilla version, FedLSA, and still relies on a heterogeneity assumption around the optimum. Another work on FSA shows the presence of a persistent bias term [59, Theorem 2], which cannot be controlled with increased local computation or even more communication rounds.
Impact of Number of Clients $M$: The sample complexity of all three algorithms scales with $1/M$, demonstrating the benefits of collaboration in FL with heterogeneous Markovian data streams and smooth non-convex objective functions.